# From chiral EFT to perturbative QCD: a Bayesian model mixing approach to symmetric nuclear matter

### v2

@@ Abstract
Constraining the equation of state (EOS) of strongly interacting, dense matter is the focus of intense experimental, observational, and theoretical effort. Chiral effective field theory ($\chi$EFT) can describe the EOS between the typical densities of nuclei and those in the outer cores of neutron stars while perturbative QCD (pQCD) can be applied to properties of deconfined quark matter, both with quantified theoretical uncertainties. However, describing the complete range of densities between nuclear saturation and an almost-free quark gas with a single EOS that has well-quantified uncertainties is a challenging problem. In this work, we argue that Bayesian multi-model inference from $\chi$EFT and pQCD can help bridge the gap between the two theories. We develop a correlated Bayesian model mixing approach that uses a Gaussian Process (GP) to assimilate different information into a single QCD EOS for symmetric nuclear matter. In the present implementation, this mixed EOS is informed solely by those of $\chi$EFT and pQCD, together with the associated truncation errors. The GP is trained on the pressure as a function of number density in the low-density and high-density regions where $\chi$EFT and pQCD are, respectively, valid. We impose priors on the GP kernel hyperparameters to suppress unphysical correlations between these regimes. This results in smooth $\chi$EFT-to-pQCD curves for both the pressure and the speed of sound. We show that using uncorrelated pointwise mixing requires uncontrolled extrapolation of at least one of $\chi$EFT or pQCD into regions where the perturbative series breaks down and leads to an acausal EOS. We also discuss future extensions and applications to neutron star matter guided by recent EOS constraints from nuclear theory, nuclear experiment, and multi-messenger astronomy.
@@ I Motivation
Our specific implementation of BMM and its application to symmetric two-flavor strongly interacting matter is described in what follows. Detailed descriptions of the $\chi$EFT and pQCD equations of state at fixed number density $n$ are presented in Sec. II, as is a determination of the truncation uncertainties in both theories. Subsection II.4 then discusses other theoretical information on the EOS that could be incorporated in our mixing. The correlated BMM approach highlighted in this work is described in Sec. III, with particular attention given to the role of GP correlation lengths. Constraining these correlation lengths according to expectations regarding the physics of the quark-hadron transition yields larger uncertainties than those obtained in the absence of such constraints. In Sec. IV, we illustrate the perils of using a pointwise BMM approach. We summarize and discuss the implications of our findings for future work in Sec. V. Appendix A discusses how we obtain the pQCD pressure as a function of the baryon density using the Kohn-Luttinger-Ward formalism Kohn and Luttinger (1960); Luttinger and Ward (1960) and Appendix B details how we compute the pQCD speed of sound at a given density consistently up to a finite order in $\alpha_{s}$. Our software implementation, including annotated Jupyter notebooks used to produce the results in this paper, will be made publicly available in a GitHub repository Semposki et al. (2024). Throughout the paper, we use natural units in which $\hbar=c=1$.
@@ II.1 Uncertainty quantification using the BUQEYE truncation error model
II.2 Low-density EOS from chiral EFT ($\chi$EFT)
@@
Throughout this work, we assume both NN and 3N LECs to be fixed to the values obtained by the developers of the nuclear potentials. Specifically, we focus on the pressure from $\chi$EFT up to N3LO (i.e., $k\leqslant 4$), obtained from the density derivative of $E(n)/A$, for the interactions with the momentum cutoff $\Lambda=500\,\text{MeV}$. We use these results from Ref. Drischler et al. (2020a), in which the BUQEYE truncation error model was used to quantify correlated EFT truncation errors and propagate them to derivative quantities such as the pressure.
@@
To construct the error bands, the scaled inverse $\chi^{2}$ priors on $\bar{c}$ and $\ell$ discussed in the previous section were used, with degrees of freedom $\nu_{0}=10$ and scale parameter $\tau_{0}^{2}=(\nu_{0}-2)/\nu_{0}$ to enforce the condition of naturalness of the known coefficients $c_{n}(x)$. A white noise kernel was added to the squared-exponential RBF kernel to account for the higher-order uncertainty from the MBPT calculations (see Ref. Drischler et al. (2020a) for more details). The 68% (i.e., $1\sigma$) intervals in Fig. 1 indicate good convergence properties at each order around and below $n_{0}$. We point out that though the N3LO band may appear wider than the N2LO band at densities $\simeq 2n_{0}$, this is not the case when comparing the size of these uncertainty bands numerically.
@@
II.3 High-density EOS from perturbative QCD (pQCD)
@@ III.2 Implementation and results
To apply the formalism of the previous subsection, we must choose the prior on the underlying theory $\kappa_{f}(x,x^{\prime})$. Here, we adopt the squared-exponential Radial Basis Function (RBF) kernel with a marginal variance $\bar{c}^{2}$. This choice of prior for the function space of $F$ carries with it several assumptions. First, $F$ is assumed to be infinitely differentiable: discontinuous phase transitions are precluded by this aspect of the prior. [^7] Note that smooth (e.g., hadron-to-quark) crossover transitions are consistent with this approach Baym et al. (2018). See, e.g., Refs. Kapusta and Welle (2021); Constantinou et al. (2021) for recent work. Second, this is a stationary kernel. An important technical point is that the GP is formulated in $\ln({n})$ space, rather than in the density $n$ itself. This is essential if pQCD is to be described by a stationary GP, since $\alpha_{s}$ runs logarithmically. Nevertheless, the assumption of stationarity means that we are implicitly assuming there is persistence in both the size and length scale of the variability in the dense matter EOS. Neither of these assumptions have to be satisfied by QCD. Testing their validity is an important topic for future investigations. But, as usual, Bayesian methods provide a clear way to delineate the consequences of such assumptions for the quantity of interest, i.e., the dense matter EOS.
@@
Table 3: A set of single values of each EOS pictured in Fig. 7. As in the main text, $\chi$EFT is taken at N3LO, and pQCD at N2LO. The mixed model is the correlated Bayesian model mixing result from the GP implementation in Sec. III.2, calculated using the two chosen lower cutoffs of the pQCD training data (Case (1): $n\geq 20n_{0}$ and Case (2): $n\geq 40n_{0}$). The chemical potential is computed using the perturbative expansion in the KLW formalism. Higher-precision, machine-readable data will be found in our GitHub repository Semposki et al. (2024).
@@
To provide easy access to our results, in Table 3, we provide the mean and standard deviation of the $P(n)$ distribution in our mixed-model EOS at densities of $4n_{0}$, $12n_{0}$, and $36n_{0}$. We list values for the EOS trained on both pQCD training cutoffs ($\geq 20n_{0}$ and $\geq 40n_{0}$). For comparison, we also provide the pQCD $P(n)$, and its uncertainty, at these same densities for our canonical choice of $X=1$. Full results will be found in the open-source GitHub repository that will accompany this work Semposki et al. (2024).
@@ IV.1 Formalism
To extrapolate the $\chi$EFT GP mean function as far as is needed to generate the result shown in Fig. 8, we employ the PAL equation of state Prakash et al. (1988); Prakash (1994), described in terms of energy per particle as [^9] To avoid confusion with other quantities in this paper, we divert in Eq. (IV.1) from the conventional notation of Ref. Prakash et al. (1988), and use $\beta$ and $\beta^{\prime}$ in place of $B$ and $B^{\prime}$, as well as $\zeta$ in place of $\sigma$.
@@ V Summary and outlook
A further and necessary assumption of any Bayesian analysis is the choice of prior on the EOS, which is chosen here to be a stationary GP. The assumption that a single stationary kernel describes the QCD EOS across all densities means that we can estimate that kernel’s length scale and marginal variance from both the pQCD and $\chi$EFT training data. The GP that describes matter in the intermediate-density region is then inconceivably Goldman well constrained if the MLE estimate of the correlation length from the data sets of $\chi$EFT and pQCD is large, as we saw in Fig. 6. To remedy this we assigned hyperpriors to the length scale and marginal variance of the mixed model GP and employed the MAP value instead of the MLE for these quantities. The hyperpriors reflect our knowledge that $\chi$EFT and pQCD uncertainties should be uncorrelated with one another in their respective dominant regions because of the transition from hadron to quark degrees of freedom. Our results for this, our preferred, analysis are shown in Fig. 7, for both chosen pQCD cutoffs of $20n_{0}$ and $40n_{0}$.
@@
The smooth, squared-exponential RBF kernel also precludes discontinuous phase transitions and other interesting structures of the types discussed in Ref. Mroczek et al. (2023). Designing advanced GP kernels to model specific EOS features, such as limited differentiability and multiple lengthscales, is an important task for future investigation. The results obtained in the model-mixed EOS with different GP kernels can then be confronted with neutron star observations and data from HIC, which has recently been used for EOS inference Huth et al. (2022); Yao et al. (2023); Komoltsev et al. (2023). Such data can guide the kernel design and selection. This process would benefit from, e.g., FRIB400 FRIB Science Community (2021); Nuclear Science Advisory Committee (2023) (NSAC), the proposed 400 MeV/u (uranium) energy upgrade of the Facility for Rare Isotope Beams (FRIB) at Michigan State University, which would provide new experimental constraints on dense matter up to twice the saturation density.
@@ Acknowledgements.
We thank Jens Braun, Tyler Gorda, Yoon Gyu Lee, Débora Mroczek, Jaki Noronha-Hostler, and John Yannotty for invaluable insights on several aspects of this work, and Pawel Danielewicz for sharing the HIC constraints on the pressure. We are also grateful to Andrius Burnelis, Andreas Ekström, Christian Forssén, Yuki Fujimoto, Pablo Giuliani, Sudhanva Lalit, Matt Pratola, Frederi Viens, and Christian Weiss for excellent discussions and advice. We acknowledge the hospitality of Chalmers University of Technology (D.R.P. and A.C.S.) and the Facility for Rare Isotope Beams (A.C.S.) during the completion of this work. C.D. thanks the BAND collaboration for their hospitality and encouragement. This research was supported by the CSSI program Award OAC-2004601 (BAND collaboration Bayesian Analysis of Nuclear Dynamics Framework project(2020) (BAND)) (A.C.S., R.J.F., D.R.P.), by the U.S. Department of Energy, Office of Science, Nuclear Physics, under Award DE-FG02-93-40756 (A.C.S., D.R.P.), by the FRIB Theory Alliance, under Award DE-SC0013617 (C.D.), by the National Science Foundation Award No. PHY-2209442 (J.A.M, R.J.F.), and by the Swedish Research Council via a Tage Erlander Professorship, Grant No. 2022-00215 (D.R.P.).
@@ B.1 Calculating P($\mu$)
indicating an approach to the conformal limit from below at the N2LO level for symmetric matter. [^10] A comprehensive Wolfram Mathematica notebook detailing this calculation will be made publicly available in the GitHub repository for this project. Semposki et al. (2024).

### v3

@@ Abstract
Constraining the equation of state (EOS) of strongly interacting, dense matter is the focus of intense experimental, observational, and theoretical effort. Chiral effective field theory ($\chi$EFT) can describe the EOS between the typical densities of nuclei and those in the outer cores of neutron stars while perturbative QCD (pQCD) can be applied to properties of deconfined quark matter, both with quantified theoretical uncertainties. However, describing the full range of densities in between with a single EOS that has well-quantified uncertainties is a challenging problem. Bayesian multi-model inference from $\chi$EFT and pQCD can help bridge the gap between the two theories. In this work, we introduce a correlated Bayesian model mixing framework that uses a Gaussian Process (GP) to assimilate different information into a single QCD EOS for symmetric nuclear matter. The present implementation uses a stationary GP to infer this mixed EOS solely from the EOSs of $\chi$EFT and pQCD while accounting for the truncation errors of each theory. The GP is trained on the pressure as a function of number density in the low- and high-density regions where $\chi$EFT and pQCD are, respectively, valid. We impose priors on the GP kernel hyperparameters to suppress unphysical correlations between these regimes. This, together with the assumption of stationarity, results in smooth $\chi$EFT-to-pQCD curves for both the pressure and the speed of sound. We show that using uncorrelated mixing requires uncontrolled extrapolation of at least one of $\chi$EFT or pQCD into regions where the perturbative series breaks down and leads to an acausal EOS. We also discuss extensions of this framework to non-stationary and less differentiable GP kernels, its future application to neutron-star matter, and the incorporation of additional constraints from nuclear theory, experiment, and multi-messenger astronomy.
@@ I Motivation
Our specific implementation of BMM and its application to symmetric two-flavor strongly interacting matter is described in what follows. Detailed descriptions of the $\chi$EFT and pQCD equations of state at fixed number density $n$ are presented in Sec. II, as is a determination of the truncation uncertainties in both theories. Subsection II.4 then discusses other theoretical information on the EOS that could be incorporated in our mixing. The correlated BMM approach highlighted in this work is described in Sec. III, with particular attention given to the role of GP correlation lengths. Constraining these correlation lengths according to expectations regarding the physics of the quark-hadron transition yields larger uncertainties than those obtained in the absence of such constraints. In Sec. IV, we illustrate the perils of using a pointwise BMM approach. We summarize and discuss the implications of our findings for future work in Sec. V. Appendix A discusses how we obtain the pQCD pressure as a function of the baryon density using the Kohn-Luttinger-Ward formalism Kohn and Luttinger (1960); Luttinger and Ward (1960) and Appendix B details how we compute the pQCD speed of sound at a given density consistently up to a finite order in $\alpha_{s}$. Our software implementation, including annotated Jupyter notebooks used to produce the results in this paper, is publicly available in a GitHub repository Semposki et al. (2024). Throughout the paper, we use natural units in which $\hbar=c=1$.
@@ II.1 Uncertainty quantification using the BUQEYE truncation error model
II.2 Low-density equation of state from chiral EFT ($\chi$EFT)
@@
Throughout this work, we assume both $NN$ and $3N$ LECs to be fixed to the values obtained by the developers of the nuclear potentials. Specifically, we focus on the pressure from $\chi$EFT up to next-to-next-to-next-to leading order (N3LO; i.e., $k\leqslant 4$), obtained from the density derivative of $E(n)/A$, for the interactions with the momentum cutoff $\Lambda=500\,\text{MeV}$. We use these results from Ref. Drischler et al. (2020a), in which the BUQEYE truncation error model was used to quantify correlated EFT truncation errors and propagate them to derivative quantities such as the pressure.
@@
To construct the error bands, the scaled inverse $\chi^{2}$ priors on $\bar{c}$ and $\ell$ discussed in the previous section were used, with degrees of freedom $\nu_{0}=10$ and scale parameter $\tau_{0}^{2}=(\nu_{0}-2)/\nu_{0}$ to enforce the condition of naturalness of the known coefficients $c_{n}(x)$. A white noise kernel was added to the squared-exponential RBF kernel to account for the higher-order uncertainty from the MBPT calculations (see Ref. Drischler et al. (2020a) for more details). The 68% (i.e., $1\sigma$) intervals in Fig. 1 indicate good convergence properties at each order around and below $n_{0}$. We point out that though the N3LO band may appear wider than the next-to-next-to leading order (N2LO) band at densities $\simeq 2n_{0}$, this is not the case when comparing the size of these uncertainty bands numerically.
@@
II.3 High-density equation of state from perturbative QCD (pQCD)
@@ III.2 Implementation and results
To apply the formalism of the previous subsection, we must choose the prior on the underlying theory $\kappa_{f}(x,x^{\prime})$. Here, we adopt the squared-exponential Radial Basis Function (RBF) kernel with a marginal variance $\bar{c}^{2}$. This choice of prior for the function space of $F$ carries with it several assumptions. First, $F$ is assumed to be infinitely differentiable. Second, this is a stationary kernel. An important technical point is that the GP is formulated in $\ln({n})$ space, rather than in the density $n$ itself. This is essential if pQCD is to be described by a stationary GP, since $\alpha_{s}$ runs logarithmically. Nevertheless, the assumption of stationarity means that we are implicitly assuming there is persistence in both the size and length scale of the variability in the dense matter EOS. Neither of these assumptions have to be satisfied by QCD. And indeed, they preclude discontinuous phase transitions. (Smooth, e.g., hadron-to-quark, crossover transitions are consistent with our assumptions Baym et al. (2018). See, e.g., Refs. Kapusta and Welle (2021); Constantinou et al. (2021) for recent work.) But, as usual, Bayesian methods provide a clear way to delineate their consequences for the quantity of interest. We are presently investigating the effect that relaxing the assumptions implemented here has on the inference of the dense matter EOS at intermediate densities Semposki et al. (2025).
@@
Table 3: A set of single values of each EOS pictured in Fig. 7. As in the main text, $\chi$EFT is taken at N3LO, and pQCD at N2LO. The mixed model is the correlated Bayesian model mixing result from the GP implementation in Sec. III.2, calculated using the two chosen lower cutoffs of the pQCD training data (Case (1): $n\geq 20n_{0}$ and Case (2): $n\geq 40n_{0}$). The chemical potential is computed using the perturbative expansion in the KLW formalism. Higher-precision, machine-readable data can be found in our GitHub repository Semposki et al. (2024).
@@
To provide easy access to our results, in Table 3, we provide the mean and standard deviation of the $P(n)$ distribution in our mixed-model EOS at densities of $4n_{0}$, $12n_{0}$, and $36n_{0}$. We list values for the EOS trained on both pQCD training cutoffs ($\geq 20n_{0}$ and $\geq 40n_{0}$). For comparison, we also provide the pQCD $P(n)$, and its uncertainty, at these same densities for our canonical choice of $X=1$. Full results are found in the open-source GitHub repository that accompanies this work Semposki et al. (2024).
@@ IV.1 Formalism
To extrapolate the $\chi$EFT GP mean function as far as is needed to generate the result shown in Fig. 8, we employ the PAL equation of state Prakash et al. (1988), described in terms of energy per particle as [^8] To avoid confusion with other quantities in this paper, we divert in Eq. (IV.1) from the conventional notation of Ref. Prakash et al. (1988), and use $\beta$ and $\beta^{\prime}$ in place of $B$ and $B^{\prime}$, as well as $\zeta$ in place of $\sigma$.
@@ V Summary and outlook
A further and necessary assumption of any Bayesian analysis is the choice of prior on the EOS, which is chosen here to be a stationary GP. The assumption that a single stationary kernel describes the QCD EOS across all densities means that we can estimate that kernel’s length scale and marginal variance from both the pQCD and $\chi$EFT training data. The GP that describes matter in the intermediate-density region is then inconceivably well constrained if the MLE estimate of the correlation length from the data sets of $\chi$EFT and pQCD is large, as we saw in Fig. 6. To remedy this we assigned hyperpriors to the length scale and marginal variance of the mixed model GP and employed the MAP value instead of the MLE for these quantities. The hyperpriors reflect our knowledge that $\chi$EFT and pQCD uncertainties should be uncorrelated with one another in their respective dominant regions because of the transition from hadron to quark degrees of freedom. Our results for this, our preferred, analysis are shown in Fig. 7, for both chosen pQCD cutoffs of $20n_{0}$ and $40n_{0}$.
@@
The smooth, squared-exponential RBF kernel also precludes discontinuous phase transitions and other interesting structures of the types discussed in Ref. Mroczek et al. (2024). Advanced GP kernels that have limited differentiability and multiple lengthscales can be used to model specific features in the EOS. Their implementation within our framework is a non-trivial extension of the studies carried out here, and will be the subject of a forthcoming paper Semposki et al. (2025). The results obtained in the model-mixed EOS with different GP kernels can then be confronted with neutron star observations and data from HIC, which has recently been used for EOS inference Huth et al. (2022); Yao et al. (2024); Komoltsev et al. (2024). Such data can guide the kernel design and selection. This process would benefit from, e.g., FRIB400 FRIB Science Community (2021); Nuclear Science Advisory Committee (2023) (NSAC), the proposed 400 MeV/u (uranium) energy upgrade of the Facility for Rare Isotope Beams (FRIB) at Michigan State University, which would provide new experimental constraints on dense matter up to twice the saturation density.
@@ Acknowledgements.
We thank Jens Braun, Tyler Gorda, Yoon Gyu Lee, Débora Mroczek, Jaki Noronha-Hostler, and John Yannotty for invaluable insights on several aspects of this work, and Pawel Danielewicz for sharing the HIC constraints on the pressure. We are also grateful to Andrius Burnelis, Andreas Ekström, Christian Forssén, Yuki Fujimoto, Pablo Giuliani, Sudhanva Lalit, Matt Pratola, Frederi Viens, and Christian Weiss for excellent discussions and advice. We acknowledge the hospitality of Chalmers University of Technology (D.R.P. and A.C.S.) and the Facility for Rare Isotope Beams (A.C.S.) during the completion of this work. C.D. thanks the BAND collaboration for their hospitality and encouragement. A.C.S., R.J.F., and C.D. are also all grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and support. This research was supported by the CSSI program Award OAC-2004601 (BAND collaboration Bayesian Analysis of Nuclear Dynamics Framework project(2020) (BAND)) (A.C.S., R.J.F., D.R.P.), by the U.S. Department of Energy, Office of Science, Nuclear Physics, under Award DE-FG02-93-40756 (A.C.S., D.R.P.), by the FRIB Theory Alliance, under Award DE-SC0013617 (C.D.), by the National Science Foundation Award Nos. PHY-2339043 (C.D.) and PHY-2209442 (J.A.M, R.J.F.), by the NUCLEI SciDAC program under award DE-FG02-96ER40963 (R.J.F.), by the Swedish Research Council via a Tage Erlander visiting Professorship, Grant No. 2022-00215 (D.R.P.), and by the US National Science Foundation via Grant PHY-2020275, N3AS PFC (D.R.P.).
@@ B.1 Calculating P($\mu$)
indicating an approach to the conformal limit from below at the N2LO level for symmetric matter. [^9] A comprehensive Wolfram Mathematica notebook detailing this calculation is publicly available in the GitHub repository for this project. Semposki et al. (2024).